OpenStudy (anonymous):

would -2x times x^2=-2x^2?

OpenStudy (anonymous):

\[-2x\times x^2=(-2\times x)\times(x\times x)=-2\times x\times x\times x=?\] :)

OpenStudy (anonymous):

solve \[-2\times x\times x\times x=\]and that will be your answer :) @jewjewbri